### v3

@@ 1 The article must be peer-reviewed and published at a conference, workshop, or journal
4 The study must discuss approaches for identifying and measuring the performance of SATD detection
5 The study must be published between the years 2014 and 2024
@@ 5 All other secondary and tertiary studies
To further enhance the comprehensiveness of the review, a bidirectional snowballing technique 30 was employed. Both forward and backward snowballing were performed, meaning that the references cited in the selected studies, as well as the studies that cited them, were examined to identify any additional relevant research that might have been overlooked in the initial search. This process resulted in the inclusion of 5 additional studies. The second author actively contributed to the snowballing process, assessing the relevance and quality of newly identified studies.
In the final stage, the 63 studies selected from the full-text filtering process were combined with the 5 studies identified through bidirectional snowballing, resulting in a total of 68 primary studies. This collaborative effort between the authors ensured that the selection process was both thorough and objective, ultimately leading to a comprehensive review of the existing literature on SATD detection approaches.
@@ 3.4 Quality Assessment Criteria
After finalizing the list of articles for the SLR, we performed a quality assessment to ensure that the selected papers provided the necessary information to answer our research questions. The quality of the selected articles was assessed using the questions presented in Table 5, following the procedure recommended by Kitchenham 26. All 68 relevant articles we discovered passed the quality evaluation, and thus, our SLR relies on these 68 primary studies.
Table 5: Quality assessment criteria
@@ 3 Does the paper describe a research methodology?
4 Does the article employ an algorithm or model for SATD detection?
@@ 5 Does the study provide research findings and answer the research questions?
6 Does the article report metric evaluation performance or at least report results of the approach?
@@ 7 Does the article clearly state the dataset used?
All of the listed articles were published between January 2014 and June 2024. Our searches for articles were completed by the end of June 2024, supplemented by articles obtained through bidirectional snowballing. As a result, it is highly unlikely that we missed any published articles during our period due to publication delays.
@@ 3.5 Data Extraction
We selected the final relevant articles after applying the inclusion and exclusion criteria and conducting the quality evaluation. We determined the required information and attributes for data summarization to address our research questions. This facilitated the data extraction process, where we read the full text of each article and extracted the relevant information based on specified attributes. To answer the research questions mentioned in Subsection 3.1, we extracted data items listed in Table 6 from each selected study. The extracted data was then entered into a spreadsheet for further analysis.
Table 6: Data extracted from each study
@@ 6 Algorithms used Algorithm or model used in the study RQ1, RQ2, RQ4
7 TD Types The types of TD, e.g., design or requirement TD R2, RQ3
8 Number of Class Number of class(es)/TD types used R2, RQ3
9 Dataset used Data source and metadata of dataset used R1, RQ4
10 Metrics Evaluation used Types of metric evaluations used for performance measurement RQ4
11 Performance Evaluations Performance measurement results using metric evaluations RQ4
@@ 4 Results
This section addresses the posed research questions. To that purpose, the outcomes of the 68 primary studies are presented and discussed.
@@
4.1 Results - RQ1: Approaches used for detecting SATD in software development
As part of the results for RQ1, we categorized the relevant studies into three main groups based on their approach: papers using NLP, those employing machine learning (ML), and studies utilizing deep learning (DL) or Transformers (see Figure 3 and Table 7). Any disagreements that arose during the grouping process were resolved through discussion.
Figure 3 shows the number of papers published by approach: NLP, ML, DL, and Transformers. Over the past decade, the identification of SATD has evolved significantly, driven by advancements in NLP and ML techniques. Early methods relied primarily on traditional NLP techniques, focusing on keyword detection and textual patterns in code comments. As depicted in Figure 3, NLP dominated initially, with limited exploration of advanced computational methods. From 2017 to 2019, ML approaches were increasingly introduced to improve SATD detection accuracy. Between 2020 and 2022, the use of DL and Transformer architectures surged, marking a shift toward more sophisticated, data-driven techniques. These models captured complex patterns in software artifacts, improving SATD detection. As of 2023 and 2024, DL and Transformer models continue to lead SATD detection efforts.
Figure 3: Articles count throughout the years based on their approaches
@@ 4.1 Results - RQ1: Approaches used for detecting SATD in software development
Table 7: Approach identified. Studies that fall into more than one approach are highlighted with square brackets
Approach Paper Reference Technique examples
NLP No. 8, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49 POS tagging, Regular Expressions, Textual patterns, etc.
Machine Learning No. 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64 65, 66, 67, 68, 69, 70, 71, 72,[73, 74, 75, 76] NB, ME, SVM, kNN, etc.
Deep Learning No. 77, 78, 79, 80, 81, 82, 82, 83, 84, 85, 86, 87, 88, 89 90, 91, 92, 93, 94,[73, 74, 75, 76] CNN, LSTM, BiLSTM, etc.
Transformers No. 95, 96, 97, 98, 99, 100,[73, 74, 75] BERT, ALBERT, GPT, etc.
NLP approaches have played a fundamental role in detecting SATD. These methods generally concentrate on text preprocessing tasks that encompass stemming, lemmatization, and stopwords removal. The objective is to process and evaluate textual data from code comments or documentation, where technical debt is frequently acknowledged by the authors themselves. As presented in Table 7, other methods include manual analysis and Parts-of-Speech (POS) tagging 36, 40, 41, 49, which aid in classifying and identifying relevant information. Regular expressions 38, 46, keyword-labeled 61, and textual patterns 45 are commonly employed to semi-automate the identification of certain technical debt indicators in large codebases.
@@
As the complexity of software systems increased, so has the demand for more sophisticated approaches in technical debt identification, resulting in the use of ML classifiers. These algorithms include Naïve Bayes (NB) 50, 55, 59, Maximum Entropy (ME) 51, 67, Support Vector Machines (SVM) 52, and k-Nearest Neighbor (kNN). Logistic Regression (LR) 101 and Random Forest (RF) 56, 43 are also employed, with various degrees of accuracy and interpretability in detecting patterns linked with technical debt. Furthermore, XGBoost 66, 68, a relatively new addition to the ML classifier, improves performance by refining gradient boosting approaches.
@@
The rise of DL has introduced even more advanced techniques for detecting SATD. Convolutional Neural Network (CNN) 77, 83, 85, 90, Long Short-Term Memory (LSTM) 79, 81, and Bidirectional Long Short-Term Memory (BiLSTM) 78, 82 are among the most prominent techniques used. They are particularly useful in processing structured data and identifying patterns that are not easily captured by traditional ML algorithms. Deep learning approaches have been extensively studied and are shown to be more critical in automatically detecting technical debt.
@@
Transformers are the most recent development in the detection of SATD. By offering excellent context understanding and language representation, these models, namely Bidirectional Encoder Representations from Transformers (BERT) 95, 96, 100, 97, A Lite BERT (ALBERT) 98, Robustly Optimized BERT (RoBERTa) 98, and DistilBERT 98, have transformed the field of NLP. Transformers excel at handling large-scale data and understanding the nuances in code comments and other textual software artifacts that signify technical debt. The literature reflects a growing trend towards using these models to achieve higher accuracy and more robust identification processes, especially in dealing with large and complex datasets typical of modern software systems.
Answer to RQ1 Four main approaches have been identified to detecting SATD: traditional NLP, ML, DL, and Transformers. Over time, the focus has shifted from basic NLP techniques to more advanced methods like ML and DL, with transformers now leading the way, due to their effectiveness in capturing complex textual patterns.
@@ 4.2 Results - RQ2: Proposed methodologies for detecting SATD in software development
To address this, we categorized the methodologies into three main types: manual, semi-automated, and automated. These categories represent the evolution in improving SATD detection’s accuracy, efficiency, and scalability. Manual methods depend on human expertise to review text artifacts, while semi-automated techniques combine human input with tools to streamline the process. Fully automated methods use advanced algorithms to detect SATD with minimal human intervention. Each approach has its own strengths and challenges, which we will explore. Disagreements in categorization were resolved through discussion.
Figure 4: Articles count throughout the years based on the methodologies
Over the past decade, methods for detecting Self-Admitted Technical Debt (SATD) have changed significantly, as shown in Figure 4. There has been a clear shift from manual to more automated techniques, driven by a growing need for efficiency and accuracy in managing technical debt in software development.
From 2014 to 2016, SATD detection mainly relied on manual analyses. This process involved developers or analysts closely examining codebases and documentation to spot technical debt. While these manual methods provided valuable insights, they were time-consuming and susceptible to human error, highlighting the demand for more efficient solutions.
@@
The period from 2020 to 2022 marked a significant move towards fully automated SATD detection methods. As depicted in the figure, there was a notable rise in the use of automated techniques, indicating the maturation of the field. These approaches employ advanced technologies such as natural language processing (NLP), machine learning (ML), and deep learning (DL) to automatically detect technical debt with minimal human input. As we move into 2023 and 2024, the trend toward automation in SATD detection continues to grow. The figure shows that automated methods have become the primary approach, outpacing manual and semi-automated techniques. This suggests that the future of SATD detection will likely focus on refining automated tools, possibly integrating more advanced AI and ML models.
@@
As presented in Table 8, the methodologies are classified into three broad categories: Manual, Semi-Automated, and Automated approaches. These categories are determined based on the extent of human intervention and the utilization of NLP, ML, DL, Transformers models.
Table 8: Methodologies identified
Methodology Paper Reference NLP ML DL TRANSFORMERS
Manual No. 8, 37, 59, 39, 102 No. 8, 37, 59, 39, 102 - - -
Semi-Automated No. 36, 38, 40, 61, 41, 44, 45, 65, 46, 47, 49, 48 No. 36, 38, 40, 41, 44, 45, 46, 47, 49, 48 No. 65, 61 - -
Automated No. 50, 51, 52, 53, 54, 55, 56, 57, 77, 58, 79, 60, 62, 42, 43, 63, 64, 80, 81, 82, 83, 74, 84, 101, 85, 86, 87, 66, 88, 67, 89, 73, 68, 69, 76, 90, 91, 92, 70, 93, 94, 75, 95, 96, 99, 100, 71, 72, 97, 98 - No. 50, 51, 52, 53, 54, 55, 56, 57, 58, 60, 61, 42, 43, 64, 84, 101, 66, 67, 73, 68, 69, 70, 71, 72 No. 77, 78, 79, 62, 63, 80, 81, 82, 83, 74, 85, 86, 87, 88, 89, 76, 90, 91, 92, 94, 96 No. 93, 75, 95, 99, 100, 97, 98
@@
Manual approaches significantly depend on human intervention, usually requiring the manual annotation of code comments or other textual software artifacts to detect SATD. Manual methods solely rely on NLP techniques. There is no application of ML, DL, or Transformers in the manual category, as these techniques are generally used for more automated processes. These methods are frequently time-consuming but offer high precision in identifying subtle or context-specific technical debt items. For example, Potdar and Shihab 8 conducted research using source code comments from several open-source projects. They manually classified comments as SATD or non-SATD based on specific comment patterns. Similarly, Maldonado and Shihab 37 employed manual labeling to construct a dataset for SATD that focused on several specific types of debt, including design debt, requirement debt, documentation debt, test debt, and defect debt. This strategy was critical in establishing baseline datasets that could subsequently be utilized to train automated classifiers 51.
@@
Semi-automated approaches combine manual review with automated tools to improve efficiency while retaining accuracy. Semi-automated methods predominantly rely on NLP, including part-of-speech (POS) tagging 36, and textual patterns 38; and a few of them use ML approaches 65, 61. However, there is no indication of the usage of Deep Learning or Transformers in these semi-automated techniques.

### v4

@@ 1 The article must be peer-reviewed and published at a conference, workshop, or journal
4 The study must discuss techniques or approaches for identifying and measuring the performance of SATD detection
5 The study must have been published between 2014 and April 2025
@@ 5 All other secondary and tertiary studies
To ensure the reliability of our inclusion and exclusion process, we followed the recommendations of Kitchenham and Charters 11, who suggest that consistency of study selection can be verified through a test-retest procedure as an alternative to inter-rater agreement. The first author conducted the initial screening of studies across all stages (metadata, research methodology, and full-text analysis). To check the consistency of the labeling process, the first author also applied a test-retest approach by re-evaluating a random sample of the studies. In addition, to minimize potential bias, a subset of 15% of the studies at each filtering stage was independently reviewed by the second author. Both validation steps yielded full agreement with the initial decisions, and no disagreements were found, indicating that the risk of misclassification was negligible. The replication package, including the screening spreadsheet and supplementary material, is publicly available.
After completing these validation steps, both authors agreed to select 69 relevant studies from six online databases for inclusion in the systematic review. This rigorous and transparent process ensured the inclusion of only the most relevant and high-quality studies.
The replication package, including the screening spreadsheet and supplementary material, is publicly available. [^3] https://github.com/edisutoyo/satd-detection-slr
3.4 Snowballing Process
To improve the completeness of the review, we employed bidirectional snowball sampling in accordance with Wohlin’s guidelines 37. In line with the practical recommendations of Brereton et al. 34, we used Google Scholar to support both backward and forward snowballing. Backward snowballing was performed by manually screening the reference lists of the included studies, whereas forward snowballing was conducted by identifying and screening papers that cited those included studies.
For backward snowballing, references within the files of the selected studies were carefully examined. Each reference was assessed based on its title and abstract to determine its relevance. When necessary, the full texts of the referenced studies were retrieved and reviewed for further evaluation. Initially, the collected references were filtered by their titles and abstracts to assess alignment with the research questions. Subsequently, the full texts of potentially relevant studies were analyzed to determine whether they met the inclusion criteria.
To ensure the reliability of the snowballing process, both authors were involved in assessing the relevance of the identified studies. The first author initially screened the references and citations, and the second author independently reviewed the candidate papers and cross-checked their eligibility. Any disagreements were discussed until a consensus was reached. Only one candidate paper required clarification at this stage, which was resolved through discussion. This indicates that disagreements were minimal and resolvable.
This rigorous process helped uncover additional studies that might have been missed during the initial search. Five additional studies were included as a result. In the final stage, the 69 studies selected through full-text screening were combined with the five studies identified through bidirectional snowballing, yielding a total of 74 primary studies. This collaborative effort ensured that the selection process was both thorough and objective, thereby leading to a comprehensive review of the existing literature on SATD detection approaches.
@@ 3.5 Quality Assessment Criteria
After finalizing the list of articles for the systematic literature review (SLR), we performed a quality assessment to ensure that the selected papers provided the necessary information to answer our research questions. We assessed the quality of the selected articles using the questions in Table 5, following the procedure recommended by Kitchenham 32. All 74 articles deemed relevant satisfied the predefined quality assessment criteria. Accordingly, our SLR is based on these 74 primary studies.
Table 5: Quality assessment criteria.
@@ 3 Does the paper describe a research methodology?
4 Does the article employ an algorithm or technique for SATD detection?
@@ 5 Does the study provide research findings and answer the research questions?
6 Does the article report metric evaluation performance, or at least report results of the approach?
@@ 7 Does the article clearly state the dataset used?
All of the listed articles were published between January 2014 and April 2025. Our searches for articles were completed by early May 2025 and were supplemented by articles identified through bidirectional snowballing. As a result, it is highly unlikely that we missed any published articles during our period due to publication delays.
@@ 3.6 Data Extraction
We selected the final relevant articles after applying the inclusion and exclusion criteria and conducting the quality evaluation. We determined the required information and attributes for data summarization to address our research questions. This facilitated data extraction, in which we read the full text of each article and extracted the relevant information based on specified attributes. To answer the research questions mentioned in Subsection 3.1, we extracted data items listed in Table 6 from each selected study. The extracted data was then entered into a spreadsheet for further analysis.
As part of the extraction, we also examined the evaluation setup adopted by each study. Since our review specifically focuses on within-project evaluations (e.g., k-fold cross-validation or random splits performed on the same dataset or project), we excluded cross-project or out-of-distribution evaluations from the performance synthesis in RQ3. This decision was made to ensure comparability of results across studies and to avoid conflating performance metrics obtained under substantially different evaluation settings.
Table 6: Data extracted from each study.
@@ 6 Algorithms used Algorithm or model used in the study RQ1, RQ2, RQ4
7 Tool proposed Name or description of any tool proposed in the study RQ1, RQ2, RQ4
8 TD types The types of TD, e.g., design or requirement debt RQ2, RQ3
9 Number of class Number of class(es)/TD types used RQ2, RQ3
10 Textual artifacts used Data source or software artifacts used RQ1, RQ4
11 Metrics evaluation used Types of metric evaluations used for performance measurement RQ3
12 Performance summary All included studies report within-project validation (e.g., k-fold CV, random splits). Cross-project results are not synthesized in this review RQ3
@@ 13 Challenges Issues or challenges related to SATD detection RQ5
To ensure the reliability of the data extraction and categorization process, we followed the guidance of Kitchenham and Charters 11. As a ‘gold-standard procedure’, they recommend that data extraction should be performed independently by at least two reviewers whenever feasible. Because full dual independent extraction for all included studies was not feasible under our resource constraints, we applied two complementary reliability checks: (i) a test–retest procedure in which the first author re-extracted and re-categorized a random sample after an interval to assess decision stability, and (ii) an independent cross-check by the second author on 15% of the extracted items. [^4] This conforms to the guidelines as suggested in section 6.4.3 of 11: ‘Alternatively, a test-retest process can be used where the researcher performs a second extraction from a random selection of primary studies to check data extraction consistency’. The shared spreadsheet used to take these decisions is available in the supplementary material at https://github.com/edisutoyo/satd-detection-slr) These checks are in line with the guidelines of Kitchenham and Charters 11, and resulted in full agreement with the initial decisions, and no disagreements were observed. For transparency and replicability, the complete dataset and categorization results are made available in our replication package.
With the methodology in place, we present our findings from the 74 selected studies. The results are organized according to our research questions, beginning with an analysis of the techniques used for SATD detection, followed by an evaluation of automation levels, effectiveness metrics, available tools, and key challenges in this domain.
@@ 4 Results
This section addresses our research questions and discusses the findings from the 74 primary studies included in this review.
@@
4.1 Results - RQ1: Techniques employed for detecting SATD in software development across textual artifacts
To classify the relevant studies, we followed a bottom-up categorization process. During data extraction, we recorded the algorithms and models used in each primary study (see Table 6) and grouped them into higher-level categories based on commonalities observed in the literature and prior surveys. This process led to four main categories (see Table 7 and Figure 3): (i) heuristic-based approaches (rule-based, keyword, or patterns matching methods), (ii) machine learning approaches (classical supervised models such as SVM, NB, and RF), (iii) deep learning approaches (neural architectures such as CNN, LSTM, BiLSTM, and GNN), and (iv) Transformer-based approaches. While Transformers are technically a specialized form of deep learning, we treat them as a distinct category because they represent a clear methodological shift: unlike recurrent models such as LSTM, Transformers rely on self-attention mechanisms 43, are more scalable, and have become the dominant paradigm in recent SATD studies. This separation follows trends observed in both NLP and SE literature 44, 45, where Transformers are increasingly treated as a distinct family of models due to their state-of-the-art performance and widespread adoption. Moreover, recent large language models (LLMs) are built on the Transformer architecture 46, further reinforcing the distinctiveness of this category.
@@
Table 7: Technique identified. Studies that fall into more than one technique are highlighted with square brackets.
Technique Paper Reference Technique examples
Heuristic No. 8, 47, 25, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60 Annotation tags, Textual patterns matching, etc.
Machine learning No. 26, 24, 61, 62, 63, 64, 65, 66, 67, 27, 68, 69, 70, 71, 72 73, 28, 74, 75, 76, 77, 78, 79, 80, 81,[82, 83, 84, 30] NB, ME, SVM, kNN, etc.
Deep learning No. 85, 86, 87, 88, 89, 90, 90, 91, 92, 93, 94, 95, 96, 97, 81 98, 99, 100, 101, 102,[82, 83, 84, 30, 31, 103, 81] CNN, LSTM, BiLSTM, etc.
Transformer No. 104, 105, 106, 107, 108, 10, 103, 109,[82, 83, 84, 31] BERT, ALBERT, ChatGPT, etc.
The categorization process followed the reliability procedures described in Subsection 3.6, which included test-retest and cross-checking by the second author. Only one disagreement was recorded and resolved through consensus, indicating that the grouping was stable and consistent.
Figure 3 presents the distribution of published papers categorized by their detection techniques, including heuristic-based methods, ML, DL, and Transformer-based approaches. Over the past decade, various approaches to SATD detection have emerged, evolving from heuristic-based methods to more advanced ML and DL techniques. Early methods relied primarily on traditional heuristic techniques, focusing on keyword detection and textual patterns in code comments. As depicted in Figure 3, heuristic methods initially dominated, with limited exploration of advanced computational methods. From 2017 to 2019, ML techniques were increasingly introduced to improve SATD detection accuracy. Between 2020 and 2022, DL and Transformer architectures surged, marking a paradigm shift toward data-driven techniques capable of capturing more complex semantic patterns. From 2023 onward, DL and Transformer models became the predominant approaches, with Transformers slightly surpassing DL in 2024 and maintaining their momentum into early 2025. This trend highlights a growing reliance on Transformer-based models for advancing SATD detection.
Figure 3: Distribution of articles by technique over the years.
Heuristic-based techniques have long played a fundamental role in detecting SATD. These approaches typically rely on lightweight, rule-driven text processing and pattern matching rather than model training. As summarized in Table 7, heuristic pipelines may also incorporate manual inspection and Parts-of-Speech (POS) tagging 47, 50, 51, 59 to support the categorization of debt-related expressions. In addition, regular expressions 48, 56, keyword-labeled patterns 69, and other textual pattern rules 55 are widely used to semi-automate the detection of recurring SATD indicators in large codebases.
@@
As software systems have grown in complexity, the need for more advanced technical debt detection methods has intensified, driving the adoption of ML classifiers to enhance accuracy and scalability. These algorithms include Naïve Bayes (NB) 26, 64, 27, Maximum Entropy (ME) 24, 74, Support Vector Machines (SVM) 61, and k-Nearest Neighbor (kNN). Logistic Regression (LR) 110 and Random Forest (RF) 65, 53 are also employed, with various degrees of accuracy and interpretability in detecting patterns linked with technical debt. Furthermore, XGBoost 28, 75 and LightGBM 75, relatively new additions to the ML classifier, improve performance by refining gradient boosting approaches.
@@
The rise of DL has introduced even more advanced techniques for detecting SATD. Convolutional Neural Network (CNN) 85, 91, 93, 98, Long Short-Term Memory (LSTM) 87, 89, and Bidirectional Long Short-Term Memory (BiLSTM) 86, 90 are among the most prominent techniques used. They are particularly useful in processing structured data and identifying patterns that are not easily captured by traditional ML algorithms. DL techniques have been extensively studied and have been shown to be more effective for automatically detecting technical debt.
@@
The transformer is the most recent development in SATD detection. By offering excellent context understanding and language representation, these models, namely Bidirectional Encoder Representations from Transformers (BERT) 104, 105, 10, 106, 31, A Lite BERT (ALBERT) 107, Robustly optimized BERT approach (RoBERTa) 107, DistilBERT 107, DistilRoBERTa 109, as well as more recent large-scale models such as Flan-T5 103 and ChatGPT 108, have transformed the field of NLP. Transformer excels at handling large-scale data and understanding the nuances in code comments and other textual software artifacts that signify technical debt. The literature reflects a growing trend toward using these models to achieve higher accuracy and more robust identification processes, particularly for large, complex datasets typical of modern software systems.
Table 8: Textual artifacts used. Studies that use more than one textual software artifact are highlighted with square brackets.
Artifact Paper Reference Heuristic ML DL Transformer
Code comment No. 8, 47, 25, 48, 24, 61, 62, 63, 64, 65, 66, 85, 67, 86, 50, 50, 68, 69, 70, 51, 52, 71, 72, 88, 89, 54, 90, 91, 55, 92, 110, 93, 94, 73, 28, 96, 74, 97, 82, 56, 75, 99, 100, 77, 101, 102, 84, 105, 108, 78, 59, 80, 103, 58, 79, 106, 81,[76, 30, 98, 10, 31] 8, 47, 25, 48, 50, 51, 54, 55, 56, 59, 58 24, 61, 62, 63, 64, 65, 66, 67, 68, 69, 52, 72, 92, 110, 73, 28, 74, 82, 75,[76], 77, 78, 79, 80 85, 86, 87, 70, 71, 88, 89, 90, 91, 94, 96, 97,[30], [98, 31], 99, 100, 102, 105, 103, 81 101, 84, 108,[10, 31], 106, 103, 81
Issue No. 26, 27, 49, 53, 111, 83, 95, 104, 109, [76, 30, 98, 10, 31] 27, 49, 111, 55 26, 92,[76] 83, 95,[30, 98, 31] 104, 109,[10, 31]
Pull request No. 107,[76, 30, 10, 31] - [76] [30, 31] 107,[10, 31]
Commit message No. 57, 60,[76, 30, 10, 31] 57, 60 [76] [30, 31] [10, 31]
Building on the approach categorization, it is essential to understand the sources or types of textual artifacts analyzed in the context of SATD. Table 8 categorizes the various studies on SATD detection based on the types of textual artifacts utilized. Code comments (such as from Java projects 24, 85, blockchain projects 59, 77, Dockerfile 55, 10, XML files 92, 10, and quantum software projects 60), issues (from issue tracking systems 30, 31 and GitHub issue reports 104, 109), pull requests, and commit messages contain valuable information for SATD detection, providing varying levels of granularity and relevance depending on the artifact analyzed. The application of heuristic, ML, DL, and Transformer-based models across these artifacts shows how approaches have evolved to handle different levels of complexity and scale.
Code comments are the most commonly analyzed artifact for SATD detection, with heuristic, ML, DL, and Transformer-based models being widely employed (e.g., studies 8, 47, 101). While issues, pull requests, and commit messages are also explored, they are less frequently studied. Early research on SATD primarily relied on heuristic-based techniques to extract patterns from developer language in code comments and issues. However, more recent studies have incorporated ML, DL, and Transformer models (e.g., 30, 101, 107, 109, 108) for a deeper and more automated analysis, especially when dealing with larger and more complex datasets. The increasing adoption of Transformer models reflects the field’s shift towards more sophisticated and scalable SATD detection across various textual artifacts.
Meanwhile, there has been a noticeable shift towards using a more diverse set of textual artifacts for detecting SATD. Furthermore, studies that utilize more than one textual software artifact (e.g., code comments combined with issue trackers or commit messages) are highlighted with square brackets. The increased use of pull requests and commit messages (e.g., 76, 30, 10, 31) as sources for SATD detection in recent studies signals a growing recognition of their value in providing temporal and contextual data on technical debt. As the field advances, there is an increasing emphasis on automation and on leveraging diverse textual artifacts to improve the scalability and reliability of SATD detection across varied software projects. The studies that employed multi-artifacts represent an effort to enhance the accuracy and robustness of SATD detection by cross-referencing different sources of developer commentary and interaction.
Our synthesis reveals a clear research bias, with most SATD detection studies focusing heavily on code comments, whereas other artifacts, such as issue trackers, commit messages, and pull requests, remain underexplored. This concentration on comments is understandable given their accessibility, yet it limits the diversity and generalizability of current approaches. Quantifying this imbalance shows that although research has advanced techniques (from heuristic to Transformer models), it has largely done so within a narrow scope of artifacts. Highlighting this gap provides guidance for future work to expand the examination of underutilized artifacts, thereby enabling a more comprehensive understanding of SATD.
Answer to RQ1 Over time, SATD detection has evolved from heuristic methods to ML and DL, with Transformer now leading due to its strength in capturing complex textual dependencies. The use of multiple artifacts has increased, yet most studies still focus on code comments, leaving issues, pull requests, and commit messages underrepresented. Quantifying this imbalance underscores the need for future research to broaden its focus toward diverse artifacts.
@@ 4.2 Results - RQ2: Proposed methodologies for detecting SATD in software development
To address this, we categorized the methodologies into three main types: manual, semi-automated, and automated. These categories represent the evolution in improving the accuracy, efficiency, and scalability of SATD detection. Manual methods depend on human expertise to review text artifacts, while semi-automated techniques combine human input with tools to streamline the process. Fully automated methods use advanced algorithms to detect SATD with minimal human intervention. Each approach has its strengths and challenges, which we will explore.
To ensure the reliability of this categorization, we followed the procedures described in Subsection 3.6, which included test-retest and cross-checking by the second author. No disagreements were recorded during this process, indicating that the grouping was consistent and reliable.
Figure 4: Distribution of articles by methodology over the years.
Over the past decade, methods for detecting SATD have changed significantly, as shown in Figure 4. There has been a clear shift from manual to more automated techniques, driven by a growing need for efficiency and accuracy in managing technical debt in software development.
From 2014 to 2016, SATD detection mainly relied on manual analyses. This process involved developers or analysts closely examining codebases and documentation to spot technical debt. While these manual methods provided valuable insights, they were time-consuming and prone to human error, underscoring the need for more efficient solutions.
@@
The period from 2020 to 2022 marked a significant shift toward fully automated SATD detection methods. As shown in the figure, there has been a notable increase in the use of automated techniques, reflecting the maturation of the field. These approaches employ advanced technologies such as ML and DL to automatically detect technical debt with minimal human intervention.
From 2023 onward, automated methods have become the dominant approach, clearly surpassing manual and semi-automated techniques. In both 2024 and early 2025, automated detection continued to expand, confirming its position as the leading methodology. This trend suggests that the future of SATD detection will likely focus on refining automated tools, with greater reliance on AI-driven techniques to enhance accuracy, efficiency, and adaptability.
@@
As presented in Table 9, the methodologies are classified into three broad categories: Manual, Semi-Automated, and Automated approaches. These categories are determined by the extent of human intervention and the use of Heuristic, ML, DL, and Transformer models.
Table 9: Methodologies identified.
Methodology Paper Reference Heuristic ML DL Transformer
Manual No. 8, 25, 27, 49, 111, 60 No. 8, 25, 27, 49, 111, 60 - - -
Semi-Automated No. 47, 48, 50, 69, 51, 54, 55, 73, 56, 57, 59, 58 No. 47, 48, 50, 51, 54, 55, 56, 57, 59, 58 No. 73, 69 - -
Automated No. 26, 24, 61, 62, 63, 64, 65, 66, 85, 67, 87, 68, 70, 52, 53, 71, 72, 88, 89, 90, 91, 83, 92, 110, 93, 94, 95, 28, 96, 74, 97, 82, 75, 76, 30, 98, 99, 100, 77, 101, 102, 84, 104, 105, 108, 10, 78, 79, 106, 107, 31, 80, 103, 81, 109 - No. 26, 24, 61, 62, 63, 64, 65, 66, 67, 68, 69, 52, 53, 72, 92, 110, 28, 74, 82, 75, 76, 77, 78, 79, 80, 103, 81 No. 85, 86, 87, 70, 71, 88, 89, 90, 91, 83, 93, 94, 95, 96, 97, 30, 98, 99, 100, 102, 105, 31, 103, 81 No. 101, 84, 104, 108, 10, 106, 107, 31, 109
@@
Manual approaches significantly depend on human intervention, typically requiring manual annotation of code comments or other textual software artifacts to detect SATD. Manual methods rely solely on heuristic-based techniques. There is no application of ML, DL, or Transformer in the manual category, as these techniques are generally used for more automated processes. These methods are frequently time-consuming but offer high precision in identifying subtle or context-specific technical debt items. For example, Potdar and Shihab 8 conducted research using source code comments from several open-source projects. They manually classified comments as SATD or non-SATD based on specific comment patterns. Similarly, Maldonado and Shihab 25 employed manual labeling to construct an SATD dataset focused on several specific types of debt, including design debt, requirement debt, documentation debt, test debt, and defect debt. This strategy was critical in establishing baseline datasets that could subsequently be utilized to train automated classifiers 24.
@@
Semi-automated approaches combine manual review with automated tools to improve efficiency while retaining accuracy. Semi-automated methods predominantly rely on heuristic-based techniques, including part-of-speech (POS) tagging 47, and textual patterns 48. While some approaches incorporate ML techniques 73, 69, there is no evidence that DL or Transformer-based models are used in these semi-automated techniques.